Find the molar mass of unknown gas which diffuses (a) twice as fast as HI (b) 0.771 times as fast as F2 (c) 1.10 times as fast as H2S
1
Expert's answer
2020-12-01T09:19:53-0500

We know the formula,

  r1/r2 = √(M2/M1)

  Where, r1 = rate of diffusion of gas with molar mass M1.

        r2 = rate of diffusion of gas with molar mass M2.


a) 

   Taking, M1 = molar mass of HI

             = 129 g/mol

   Molar mass of unknown gas = M2 g/mol

Given,

   the gas diffuses twice as fast as HI

So, 

 diffusion rate of HI, r1 ×2 = diffusion rate of unknown gas, r2

 Or, r1×2 = r2

 Or, r1/r2 = (1/2)

  

     r1/r2 = √(M2/M1)

  Or, (1/2) = √(M2/129)

  Or, 1/4 = M2/129

  Or, M2 = 129/4

        = 32.25 g/mol

So, molar mass of the unknown gas = 32.25 g/mol



b)

 Taking, M1' = molar mass of F2

              = 38 g/mol

  Molar mass of unknown gas = M2' g/mol Given,

  the gas diffuses 0.771 times faster than F2

 So, 

 diffusion rate of HI, r1 ×0.771= diffusion rate of unknown gas, r2

        Or, r1×0.771 = r2

        Or, r1/r2 = (1/0.771)

         

     r1/r2 = √(M2/M1)

       Or, (1/0.771) = √(M2/38)

       Or, 1/(0.771)² = M2/38

       Or, M2 =38/ 0.594

             = 63.97 g/mol

 So, molar mass of the unknown gas = 63.97 g/mol



C)

Taking, M1 = molar mass of H2S

             = 34 g/mol

Molar mass of unknown gas = M2 g/mol

 Given,

the gas diffuses twice as fast as H2S

 So, 

 diffusion rate of H2S, r1 ×1.10 = diffusion rate of unknown gas, r2

      Or, r1×2 = r2

      Or, r1/r2 = (1/2)

     r1/r2 = √(M2/M1)

       Or, (1/1.10) = √(M2/34)

       Or, 1/1.21= M2/34

       Or, M2 = 34/1.21

              = 28.1 g/mol

So, molar mass of the unknown gas = 28.1 g/mol